### v1

# Towards Interpretable Deep Neural Networks for Tabular Data

Abstract
Tabular data is the foundation of many applications in fields such as finance and healthcare. Although DNNs tailored for tabular data achieve competitive predictive performance, they are blackboxes with little interpretability. We introduce XNNTab, a neural architecture that uses a sparse autoencoder (SAE) to learn a dictionary of monosemantic features within the latent space used for prediction. Using an automated method, we assign human-interpretable semantics to these features. This allows us to represent predictions as linear combinations of semantically meaningful components. Empirical evaluations demonstrate that XNNTab attains performance on par with or exceeding that of state-of-the-art, black-box neural models and classical machine learning approaches while being fully interpretable.
1 Introduction
Tabular data is the most common type of data in a wide range of industries, including advertising, finance and healthcare. While deep neural networks (DNNs) achieved major advances in computer vision, their performance on tabular data remained below that of Gradient Boosted Decision Trees (GBDTs) (Chen and Guestrin, 2016; Khan et al., 2022). To improve the performance of DNNs on tabular data, several deep learning methods have been developed with specialized architectures that take into account the unique traits of tabular data (mixture of categorical, ordinal and continuous features) in order to be on par (Huang et al., 2020) or even outperform GBDTs (Chen et al., 2024). While these methods addressed the limitations of DNNs in terms of performance, DNNs still remain blackboxes, whose decisions are hard to communicate to end users (Schwalbe and Finzel, 2023).
In this paper, we address the interpretability limitation of existing DNNs and present XNNTab, a deep neural network with a sparse autoencoder (SAE) component. The SAE learns a dictionary of monosemantic, sparse features that are used for outcome prediction. We take inspiration from the success of SAEs in learning interpretable features from LLM’s internal activations (Huben et al., 2023; Yun et al., 2021). We use an automated method to assign human-understandable semantics to the learned monosemantic dictionary features. Our method is illustrated in Figure 1 and described in Section 3. We show that XNNTab outperforms both interpretable and classical blackbox models and has similar performance to GBDTs. Additionally, XNNTab is fully interpretable: the final prediction is a simple linear combination of the interpretable dictionary features, which have easily accessible human-understandable semantics.
2 Related Work
Gradient boosted decision trees (GBDTs) (Chen and Guestrin, 2016; Khan et al., 2022) have shown state-of-the art performance on tabular data (Grinsztajn et al., 2022). To close this performance gap to GBDTs, neural architectures specifically designed for tabular data have been proposed, such as attention-based architectures (Gorishniy et al., 2021; Somepalli et al., 2022; Yan et al., 2023), retrieval-augmented models (Somepalli et al., 2022; Gorishniy et al., 2023). Additional techniques include tailored regularizations (Jeffares et al., 2023), deep ensembles (Gorishniy et al., 2024), attentive feature selection strategies (Arik and Pfister, 2021), categorical feature embedding (Huang et al., 2020) and reconstruction of binned feature indices (Lee et al., 2024).
Recent methods focus not only on performance, but also aim to include interpretability. ProtoGate (Jiang et al., 2024) is an ante-hoc interpretable architecture for tabular data, which learns prototypes of classes and predicts new instances based on their similarity to class prototypes. InterpreTabNet (Si et al., 2024), a variant of TabNet (Arik and Pfister, 2021) adds post-hoc explanations by learning sparse feature attribution masks and using LLMs to interpret the learned features from the masks. Similar to ProtoGate and different to InterpetTabNet, XNNTab is intrinsically interpretable, i.e., it uses interpretable features directly for prediction. While ProtoGate relies on class prototypes, XNNTab learns relevant feature combinations that describe parts of instances, similar to part-prototype models (Elhadri et al., 2025).
3 Method
XNNTab consists of a standard neural blackbox (e.g., an MLP) to learn nonlinear latent features, and a sparse autoencoder (SAE) (Huben et al., 2023) to decompose this latent polysemantic feature space into monosemantic features (cf. Figure 1, A). The monosemantic features are in a latent space with unknown (but unique) semantics. To assign meaning to each feature, we learn rules that describe the subset of the training samples that highly activate each feature (cf. Figure 1, B). Finally, we combine the last linear layers into a single linear layer by multiplying their weight matrices (cf. Figure 1, C).
Figure 1: Architecture overview. A. An MLP learns nonlinear features, which are decomposed into monosemantic dictionary features using an SAE. B. Dictionary features are assigned a human-interpretable meaning by learning rules for the subset of training instances that highly activate a specific feature. C. Predictions are linear combinations of monosemantic features by combining the linear model components (the decoder, $M^{T}$, and the linear layer of the MLP, $W$).
3.1 Architecture
The base blackbox model is an MLP and with multiple hidden layers and a softmax output. We denote the penultimate layer as $l$, and the representations of this layer as $h_{l}$. The MLP learns a function $\hat{y}=f(x)$. $f(x)$ can be decomposed into a function $g$ learning the hidden representations $h_{l}$ and a linear model $h$ predicting the target based on $h_{l}$, i.e., $\hat{y}=h(h_{l})=h(g(x))$. $g(\cdot)$ is characterized by the parameters $\theta_{g}$, and $h(\cdot)$ is characterized by the weight matrix $W$.
For the SAE, we use the architecture introduced by Huben et al. (2023). The SAE consists of an input layer, a hidden layer and an output layer. The size of the hidden layer is $d_{hid}=R\cdot d_{in}$ with $d_{in}$ being the dimension of the MLP’s layer $l$, and $R$ a hyperparameter controlling the size of the feature dictionary. The input to the SAE are the hidden representations of the blackbox MLP $h_{l}$. The weights of the encoder and decoder are tied, and given by the matrix $M$ ($M^{T}$ for the decoder)  [^1] Additionally, the encoder has a bias vector $b$.. The output of the SAE is given by $\hat{h}_{l}=M^{T}\Bigl{(}RELU(Mh_{l}+b)\Bigr{)}$. By design, $\hat{h}_{l}$ is a sparse linear combinations of (latent) dictionary features.
3.2 Training
XNNTab is trained in four steps: i) training of the MLP’s representation, ii) training of the SAE, and iii) finetuning of the decision layer, iv) combining linear components.
Step 1: MLP training. The MLP’s parameters $\theta_{\text{MLP}}$= ($\theta_{g}$, $W$) are trained using cross-entropy loss and L1 regularization. The SAE is not present in this step.
Step 2: SAE Training. To learn a dictionary of sparse features we follow the training procedure by Huben et al. (2023). The SAE is trained to reconstruct the hidden activations $h_{l}$ of training samples $x$. The SAE uses a reconstructing loss and a sparsity loss on its hidden activations $h_{SAE}$ to learn sparse dictionary features.
Step 3: Finetuning the decision layer. We freeze the parameters $\theta_{g}$ of the MLP and $\theta_{SAE}$ of the SAE and finetune the weights $W$ in the decision layer with the same loss as in step 1 to make predictions from the reconstructions learned by the SAE $\hat{h}_{l}$.
[…]
3.3 Learning Representation Semantics
The hidden representations of the SAE represent monosemantic dictionary features in a latent space, but their semantics are unknown. To assign human-understandable semantics to those features, we follow a similar idea as in (Huben et al., 2023) for tabular data instead of text (cf. Figure 1, B). For each dimension in the latent representation of the SAE $j\in\{1,...d_{hid}\}$, we identify the subset of training samples $T_{j}$ which highly active this feature, i.e., whose activations are above a threshold $t$. We then use a rule-based classifier to learn simple decision rules describing this subset.
4 Experiments
Datasets and Evaluation Metrics. We evaluate on two benchmark datasets for structured data, Adult (ADULT(Vanschoren et al., 2014) and Churn Modelling (CHURN) [^3] https://www.kaggle.com/datasets/shrutimechlearn/churn-modelling. For ADULT, we only used the numerical features resulting in 6 features describing approx. 50k instances. CHURN contains 14 features and 10k instances. In both cases, the task is binary classification. For evaluation, we use 5-fold-cross-validation. The training/validation/test proportion of the datasets for each split are 65%/15%/20%. We report accuracy and macro F1 on the test set, averaged across all folds.
Models and Parameters. We used neural architecture search for the MLP and SAE components of XNNTab. The base MLP on ADULT has 3 layers with (97, 30, 7) neurons, the SAE $3\times 7=21$ (R=3) neurons in the hidden layer. For CHURN, the base MLP has 2 layers with (63, 24) neurons, the SAE $2\times 24=48$ (R=2) neurons in the hidden layer. See Appendix A.2 for details on all hyperparameters.
Representation Semantics. To assign human-understandable semantics to the dictionary features, we employ Skope-rules [^4] https://github.com/scikit-learn-contrib/skope-rules. Accessed July 2025. For each dictionary feature $f_{j}$, we retrieve a set $T_{j}$ of instances for which the activation of neuron $j$ is above the threshold $t=0.9$. We assign $label=1$ to instances in $T_{j}$ and $label=0$ to the remaining instances. The Skope-rules classifier learns a set of decision rules describing instances of the positive class. For each $T_{j}$, we keep the rule with the highest coverage (recall).
Results. Table 1 shows that, on both datasets, XNNTab outperforms interpretable models. On ADULT, XNNTab outperforms MLP and Random Forest and is comparable to XGBoost, on CHURN, XNNTab outperforms all baseline blackbox models including XGBoost. These results show that our model introduces interpretability to DNNs with very little compromise on performance. Rules extracted for the dictionary features on ADULT are shown in Table 2, for CHURN in Appendix B.2. The learned rules provide a global explanation of the model’s behavior (Figure 2 left), since the predictions are simply a linear combination of the learned features (rules). Figure 2, right shows that on average 8 out of the 21 features that we learn for ADULT are active. That means, one local explanation requires us to inspect 8 rules to understand the model’s prediction. However, the rules are simple to understand (two conditions long on average). The number of active features on average increase for CHURN but not the complexity of the rules (average rule length remains 2).
Table 1: Performance of blackbox (✕) and fully interpretable (✓) models on two benchmark datasets. Best values are marked bold, best values for interpretable models bold italic.
ADULT CHURN
F1-Macro Acc F1-Macro Acc
Random Forest 0.717 $\pm$ 0.003 0.806 $\pm$ 0.002 0.738 $\pm$ 0.002 0.857 $\pm$ 0.001
XGBoost 0.750 $\pm$ 0.006 0.843 $\pm$ 0.003 0.730 $\pm$ 0.000 0.854 $\pm$ 0.000
MLP 0.730 $\pm$ 0.004 0.824 $\pm$ 0.002 0.745 $\pm$ 0.001 0.857 $\pm$ 0.001
Logistic Regression 0.691 $\pm$ 0.000 0.815 $\pm$ 0.002 0.601 $\pm$ 0.006 0.808 $\pm$ 0.000
Decision Tree 0.709 $\pm$ 0.000 0.804 $\pm$ 0.000 0.718 $\pm$ 0.011 0.836 $\pm$ 0.009
XNNTab (ours) 0.733 $\pm$ 0.006 0.823 $\pm$ 0.002 0.759 $\pm$ 0.001 0.861 $\pm$ 0.005
Table 2: Dictionary features for the ADULT dataset. $|T_{j}|$ - size of the training subset that strongly activate feature $j$. Coverage of the rule reported as number of samples and percentage of samples. Table sorted by $|T_{j}|$.
j $|T_{j}|$ Description Coverage
5 6755 age $\leq 26$ and edu_num $\leq 12$ 4940 / 0.73
11 6217 age $\leq 26$ and cg $\leq 2969$ and edu_num $\leq 11$ 4722 / 0.76
16 3863 age $\leq 22$ and cg $\leq 3389$ and hpw $\leq 43$ 2604 / 0.67
15 3328 age $\leq$ and hpw $\leq$ 30.5 2082 / 0.62
9 2872 age $\leq 24$ and hpw $\leq 31$ 1919 / 0.66
21 1123 cg > 8296 and hpw > 22 751 / 0.68
6 789 age > 35 and cg > 9474 700 / 0.79
20 660 age $\leq 21$ and edu_num $\leq 9$ and cg $\leq 60025$ and hpw $\leq 23$ 406 / 0.61
1 210 cg > 22587.5 210 / 1.00
8 210 cg > 22588 and cl $\leq 78$ 210 / 1.00
13 203 age > 19 and cg > 25180 202 / 0.99
14 155 cg > 70655 and cl $\leq 78$ 155 / 1.00
ADULT CHURN
Rule Length
     min 1 1
     avg 2.2 1.8
     max 4 2
Active Features
     min 3 16
     avg 8 21.5
     max 11 28
Figure 2: Left: Decision weight matrix $W^{\prime}$ for ADULT on interpretable features. Right: Statistics on complexity of features on ADULT and CHURN.
Appendix A Implementation Details
A.1 Datasets
Table 3: Dataset overview
# Instances # Features Task
Adult 48 842 6 Binary classification
Churn Modeling 10 000 14 Binary classification
Data Details. ADULT contains demographic information such as age, education, and capital gain. The target variable is wether a person’s income is >50k or $\leq 50$. CHURN contains details of a bank’s customers, the target variable is whether the customer left the bank or continues being a customer.
Data preprocessing. We normalize the input features of ADULT. For CHURN, we drop the feature columns “RowNumber”, “CustomerId”, and “Surname”, encode the categorical columns and normalize the numerical ones to $[0,1]$ by min-max normalization.
A.2 Parameter settings
MLP. We use the Optuna library  [^5] https://optuna.org for neural architecture search and tune on the number of layers in the range of [1, 4], the dimension of the hidden layers (latent_dim) in the range of [1, 200], the learning rate ($lr$) in the range of [1e21e-2, 5e35e-3], and dropout rate ($Dropout$) in the range of [0.0, 0.5]. The parameter settings for ADULT are latent_dim = [97, 30, 7], Dropout = [0.3, 0.35, 0.23], lr=4e3lr=4e-3, and $\lambda=1e-4$ for L1 regularization. The parameter settings for CHURN are latent_dim = [63, 24], Dropout = [0.4, 0.3], lr=1e2lr=1e-2, and $\lambda=4e-4$ for L1 regularization.
XNNTab Follows the same setting details as MLP for the neural network component, the parameter settings for the sparse autoencoder are the $l1$ coefficient $\alpha=1e-3$ for ADULT and CHURN.
For the following baseline models we use the default parameter settings from scikit-learn (v1.7.1) [^6] https://scikit-learn.org/stable/
Random Forest. We tuned on number of estimators (n_estimators)(n\_estimators) from the list [100, 200].
XGBoost. We tuned on number of boosting iterations (num_boost_round)(num\_boost\_round) in the range of [100, 150, 200] and set the early stopping early_stopping_roundsearly\_stopping\_rounds to 12.
Decision Trees. We tuned on maximum depth (max_depth)(max\_depth) in the range of[5, 10, 15, 20].
Logistic Regression. We tuned on maximum of iterations (max_iter)(max\_iter) in the range of [100, 200].
For our rule-based classifier:
Skope-rules. we set precision_min=1precision\_min=1, recall_min=0.2recall\_min=0.2, and max_depth=4max\_depth=4, and the default settings for the rest of the parameters.
Appendix B Qualitative Results
B.1 ADULT
Table 4: Local explanation of an instance with positive prediction (> 50k) from ADULT. $c_{j}$ are the nonzero activations of dictionary features, the contribution of all activated features to the two output classes is $c^{\top}W^{\prime}$, where $W^{\prime}$ is the combined weight matrix.
Dataset features age fnlwgt edu_num cg cl hpw
Raw instance 48 175622 15 99999 0 60
Dictionary feature
$j$ rule $c_{j}$ $W^{\prime}_{j,0}$ $c_{j}\cdot W^{\prime}_{j,0}$ $W^{\prime}_{j,1}$ $c_{j}\cdot W^{\prime}_{j,1}$
1 cg > 22587.5 4.8735 -0.6028 -2.939 0.4310 2.101
6 age > 35 $\wedge$ cg > 9474 12.9912 -0.1803 -2.342 0.2962 3.848
8 cg > 22588 $\wedge$ cl $\leq 78$ 4.5567 0.1452 0.662 0.3139 1.431
9 age $\leq 24$ $\wedge$ hpw $\leq 31$ 5.1609 0.2681 1.384 -0.0763 -0.394
13 age > 19 $\wedge$ cg > 25180 3.9865 -0.6343 -2.528 0.4284 1.709
14 cg > 70655 $\wedge$ cl $\leq 78$ 1.8393 -0.1363 -0.251 -0.2378 -0.437
16 age $\leq 22$ $\wedge$ cg $\leq 3389$ $\wedge$ hpw $\leq 43$ 5.6880 0.1648 0.937 0.0393 0.224
20 age $\leq 21$ $\wedge$ edu_num $\leq 9$ $\wedge$ cg $\leq 60025$ $\wedge$ hpw $\leq 23$ 2.0693 0.1102 0.228 0.0218 0.045
21 cg > 8296 $\wedge$ hpw > 22 15.8202 -0.4210 -6.660 0.6348 10.039
Output Logits -11.509 18.567
Class $\mathbf{\leq 50K}$ $\mathbf{\geq 50K}$
Table 5: Local explanation of one instance with negative prediction ($\leq$ 50k) from ADULT. $c_{j}$ are the nonzero activations of dictionary features, the contribution of all activated features to the two output classes is $c^{\top}W^{\prime}$, where $W^{\prime}$ is the combined weight matrix. Dictionary features 7, 10 and 19 never exceeded the activation threshold $t=0.9$ and hence have no rules associated.
Dataset features Age fnlwgt edu_num cg cl hpw
Raw instance 18 225859 10 2907 0 30
Dictionary feature
$j$ rule $c_{j}$ $W^{\prime}_{j,0}$ $c_{j}\cdot W^{\prime}_{j,0}$ $W^{\prime}_{j,1}$ $c_{j}\cdot W^{\prime}_{j,1}$
5 age $\leq 26$ $\wedge$ edu_num $\leq 12$ 1.9102 0.5432 $\mathbf{1.0376}$ -0.6966 $\mathbf{-1.3306}$
6 age > 35 $\wedge$ cg > 9474 0.4034 -0.1803 $\mathbf{-0.0727}$ 0.2962 $\mathbf{0.1195}$
7 0.2820 0.1042 $\mathbf{0.0294}$ -0.5845 $\mathbf{-0.1648}$
9 age $\leq 24$ $\wedge$ hpw $\leq 31$ 0.9971 0.2681 $\mathbf{0.2673}$ -0.0763 $\mathbf{-0.0761}$
10 0.3919 0.3814 $\mathbf{0.1495}$ -0.3506 $\mathbf{-0.1374}$
11 age $\leq 26$ $\wedge$ cg $\leq 2969$ $\wedge$ edu_num $\leq 11$ 1.7615 0.3644 $\mathbf{0.6419}$ -0.5136 $\mathbf{-0.9047}$
14 cg > 70655 $\wedge$ cl $\leq 78$ 0.1123 -0.1363 $\mathbf{-0.0153}$ -0.2378 $\mathbf{-0.0267}$
15 age $\leq$ $\wedge$ hpw $\leq$ 30.5 1.1112 0.0624 $\mathbf{0.0693}$ -0.5594 $\mathbf{-0.6216}$
16 age $\leq 22$ $\wedge$ cg $\leq 3389$ $\wedge$ hpw $\leq 43$ 1.1850 0.1648 $\mathbf{0.1953}$ 0.0393 $\mathbf{0.0466}$
19 0.3856 0.3426 $\mathbf{0.1321}$ -0.0268 $\mathbf{-0.0103}$
20 age $\leq 21$ $\wedge$ edu_num $\leq 9$ $\wedge$ cg $\leq 60025$ $\wedge$ hpw $\leq 23$ 0.5693 0.1102 $\mathbf{0.0627}$ 0.0218 $\mathbf{0.0124}$
Output Logits 2.4662 -3.0938
Class $\mathbf{\leq 50K}$ $\mathbf{\geq 50K}$
B.2 CHURN
Table 6: Dictionary features for the CHURN dataset. $|T_{j}|$ - size of the training subset that strongly activate feature $j$. Coverage of the rule reported as number of samples and percentage of samples. Table sorted by $|T_{j}|$.
$j$ $|T_{j}|$ Description Coverage
6 4955 Age < 45 and NumOfProducts < 3 4560 / 0.91
47 4631 Age $\leq$ 55.5 and NumOfProducts $\geq$ 2 2898 / 0.62
2 4628 CreditScore > 379.0 and Age < 40 3561 / 0.76
3 4457 CreditScore > 409.0 and NumOfProducts $\geq$ 2 2939 /0.66
30 4419 Age < 45 and Balance > 76177.5 2594 /0.57
43 3966 Age $\leq$ 42 2238 / 0.62
9 2999 Age < 46.5 and NumOfProducts $\geq$ 2 2533 / 0.84
12 2574 Age < 35 and NumOfProducts $\geq$ 2 1193 / 0.46
18 1805 Geography_Spain == False and Age $\geq$ 45 1081 / 0.59
36 425 Age > 59.5 and IsActiveMember == 1 255 / 0.60
7 424 NumOfProducts $\geq$ 3 196 / 0.46
24 363 CreditScore > 361.5 and NumOfProducts $\geq$ 3 211 / 0.57
1 350 CreditScore > 379.0 and NumOfProducts $\geq$ 3 205 / 0.59
38 303 Age > 64.0 and IsActiveMember == 1 172 / 0.57
28 270 Balance $\leq$ 244642.5 and NumOfProducts $\geq$ 3 215 /0.79
40 226 NumOfProducts $\geq$ 3 196 / 0.87
11 218 Age > 62.5 and IsActiveMember == True 164 /0.74
35 150 Age $\geq$ 38 and Balance $\geq$ 75720 98 /0.65
18 146 Geography_Germany == True and IsActiveMember == False 54 /0.29
4 127 Age > 61.5 and NumOfProducts < 2 and IsActiveMember == True 101 / 0.79
16 112 Age > 66.0 and NumOfProducts < 2 and IsActiveMember == True 60 / 0.55
5 53 Age > 77 12 / 0.23